A major league baseball player finished the season with 122 singles. Five percent of his hits were triples. Two out of every fifteen hits were home runs, and he hit one more double than he did home runs. How many hits altogether did he have that season?
Answer by stanbon(75887)

A major league baseball player finished the season with 122 singles.
Five percent of his hits were triples.
Two out of every fifteen hits were home runs,
and he hit one more double than he did home runs.
-----------
How many hits altogether did he have that season?
---------------------------------------------
Let "x" be the total number of hits.
singles : 122
triples: 0.05x
home runs: (2/15)x
doubles: (2/15)x + 1
===========================
Equation:
(0.95)x = 122 + (2/15)x + (2/15)x + 1
---
0.95x = 123 + (4/15)x
---
0.95x - (4/15)x = 123
(41/60)x = 123
x = 180 hits (total # of hits)
====================================
Cheers,
Stan H.